# Protecting Your Voice: Temporal-aware Robust Watermarking

### v1

@@ Abstract
Voice watermarking, Temporal-aware watermarking, Proactive forensic
@@ I Introduction
In the field of audio and voice watermarking, it has seemingly become an established principle that achieving good robustness necessitates a frequency domain transformation. As depicted in the upper branch of Fig. 1, this holds true both in the earlier era of handcrafted watermarking [14, 15, 16, 17, 18, 19] and in the current epoch dominated by deep-learning-based approaches [20, 21, 22, 23, 24, 25, 26, 27]. Concretely, Liu et al. [23] embedded the watermark into the approximate coefficients derived from the Discrete Wavelet Transform (DWT). Fully leveraging the short time window properties of the Short-Time Fourier Transform (STFT), Liu et al. [22] employed frequency features for robust watermarking. Additionally, the advantageous properties of STFT magnitudes have motivated subsequent works, including Chen et al. [25], Pavlovic et al. [26], and O’Reilly et al. [27], to embed watermarks into these magnitudes. For DL-based approaches mentioned above, another notable phenomenon is observed: achieving the high-level robustness performance necessitates the use of an attack simulator (AS). This observation prompts us to rethink: can robustness rely solely on the attack simulator? If so, it may be possible to shift the watermark embedding process from the frequency domain to the temporal domain, potentially achieving improved timbre quality.
Figure 1: Two branches of watermarking methods. Upper: Frequency-domain watermarking requires fundamental transformations. Lower: The proposed time-domain watermarking embeds the watermark into temporal features.
@@
  • Sound Performance. Comprehensive experiments and comparisons with state-of-the-art methods demonstrate the superior fidelity and enhanced robustness.
@@ II-B Extracting Module
Architectures: Gating mechanism[29] has demonstrated its effectiveness in feature extraction for language processing tasks[30] and classification tasks[31]. To fully exploit this advantage, we have designed the gating mechanism to capture temporal domain features, resulting in a Temporal-aware Gated Convolutional Network (TGC), which incorporates the ConvBlock in extracting module. The TGC follows a dual-brach structure, where the main branch consists of a sequential arrangement of a Conv1d layer, batch normalization, and a sigmoid activation to implement the gating mechanism. In contrast, the shortcut branch contains only a Conv1d layer for acquiring fine-grained features from the input. The outputs of these two branches are merged using the Hadamard Product to produce the final output. This efficient design contributes to the lightweight nature of the ConvBlock.
@@ III Experimental Results and Analysis
This section presents experiments conducted to evaluate the proposed True method based on three primary assessment criteria: fidelity, capacity, and robustness. Additionally, comprehensive experiments and analyses are provided to compare its performance against state-of-the-art (SOTA) methods, further demonstrating its effectiveness.
@@ III-A Experimental Setup
Datasets and Baseline: In the context of voice datasets, other than conventional speech voice datasets like LJSpeech [34], LibriTTS [35], and Aishell3 [36], singing voice datasets like M4Singer [37] and Opencpop [38] are also considered. Moreover, comprehensive comparisons against SOTA watermarking methods are conducted, including handcrafted methods such as FSVC [16], Normspace [17], and PBML [19], as well as deep learning-based methods like AudioSeal [21], WavMark [25], TAWM [22], and Groot [33].
@@
Models & Training Settings: In the downsampling block of ICDE, Conv1d layers use a kernel size of 3, a stride of 2, and padding of 2, except for the last layer, which has padding of 1. In the upsampling block, Conv1d layers maintain a kernel size of 3, a stride of 1, and padding of 1, while transposed Conv1d layers have a kernel size of 3, a stride of 2, padding of 2, and output padding of 1, except for the first layer, which uses padding of 1. For the TGCs, both Conv1d layers in the two branches have a kernel size of 3, a stride of 2, and padding of 1. During training, AdamW optimizer [42] is employed with a learning rate of 2e-4. The training is conducted for 40 epochs with a batch size of 16.
@@ III-C Fidelity and Capacity
Fidelity is an essential criterion for assessing the performance of watermarking methods. Table II presents the speech and singing voice quality and watermark extraction accuracy of the proposed True. Moreover, it also provides a comparison of fidelity with SOTA methods on LJSpeech. All data are obtained by comparing the watermarked waveform with the natural waveform. From these results, it can be observed that the proposed True demonstrated excellent speech quality under reasonable watermark extraction accuracy on speech voice datasets. On the other hand, it also achieved notable fidelity and extraction accuracy on singing voice datasets. Furthermore, in comparative experiments with SOTA methods, our True ranked first in both PESQ and SSIM metrics, showcasing better speech quality.
Capacity reflects the watermark lengths accommodated by the proposed method, which is equally significant as fidelity. Table III illustrates the effectiveness of the proposed method on the LJSpeech and Opencpop datasets under various watermark capacities. The experimental results indicate that True is well-suited for high-capacity scenarios up to 500 bps. For the LJSpeech, our method achieves an average watermark extraction accuracy of 97.84% while maintaining a high fidelity with an average PESQ score of 4.6140. Similarly, the Opencpop dataset maintains an average PESQ of 4.5222 and achieves an average recovery accuracy of 98.26%. When capacity exceeds 600 bps, it would lead to a significant decrease in criteria.
@@ III-D Robustness
We evaluated the robustness of various datasets at a capacity of 100 bps. GN, Pink noise (PN), LP, BP, HP, SPS, ReS, Echo, TS, and Dither were employed for validation. Table IV illustrates the watermark extraction accuracy under these attacks for each individual dataset. The results indicate that the proposed method exhibits considerable robustness for singing voice. Concretely, on the singing voice datasets, it achieves watermark extraction accuracies of 99.61% and 97.82% under the noise level of 20 dB. Furthermore, when subjected to remanent attacks, True achieves average recovery accuracies of 99.45% and 96.85%, respectively. Although the performance on the speech voice is somewhat diminished compared to the signing voice datasets, it still exhibits desirable robustness, achieving average accuracies of 92.33%, 94.85%, and 91.91% across the singing voice datasets.
The proposed True was further compared with SOTA methods on the LJSpeech dataset, assessing its robustness across multiple voice post-processing operations, including GN, PN, BP, SPS, and Echo. The extraction accuracies of various methods are illustrated in Table V. To ensure a fair comparison, we conducted two distinct sets of experiments with capacities of 32 bps and 100 bps, respectively. At a capacity of 32 bps, our proposed method demonstrates a higher watermark extraction accuracy compared to other SOTA methods. Notably, True demonstrates remarkable resilience, maintaining an average extraction accuracy of 99.14% even after subjecting the waveform to six distinct attacks. Our True maintained strong robustness in the 100 bps experimental configuration, particularly in handling PN and SPS. While the accuracy under a noise level of 10 dB was marginally lower compared to the Groot, True still significantly exceeded the other two methods. Although not achieving the highest accuracy in addressing other attacks, True manifested substantial robustness.
@@ IV Conclusion
In this study, we propose a temporal-aware robust watermarking named True, designed to proactively protect copyrights of multi-type waveforms (speech voice and singing voice). To this end, an integrated content-driven encoder is meticulously designed to facilitate the seamless incorporation of the watermark with temporal representations of the waveform and to reconstruct the watermarked waveform directly. During extraction, the temporal-aware gate convolutional network is further developed to capture fine-grained features of the watermarked waveform after attack simulation, boosting the accuracy of watermark recovery. The proposed True, outperforms baseline methods in terms of fidelity and can handle large capacity of up to 500 bps. In addition, it demonstrates exceptional robustness against a wide range of common waveform attacks.

### v2

@@ Abstract
Audio watermarking, Temporal-aware watermarking, Proactive forensics
@@ I Introduction
In the field of audio and voice watermarking, it has seemingly become an established principle that achieving good robustness necessitates a frequency domain transformation (FDT). As depicted in the upper branch of Fig. 1, this holds true both in the earlier era of handcrafted watermarking [14, 15, 16, 17, 18, 19] and in the current epoch dominated by deep-learning-based approaches [20, 21, 22, 23, 24, 25, 26, 27]. Concretely, Liu et al. [23] embedded the watermark into the approximate coefficients derived from the Discrete Wavelet Transform (DWT). Fully leveraging the short time window properties of the Short-Time Fourier Transform (STFT), Liu et al. [22] employed frequency features for robust watermarking. Additionally, the advantageous properties of STFT magnitudes have motivated subsequent works, including Chen et al. [25], Pavlovic et al. [26], and O’Reilly et al. [27], to embed watermarks into these magnitudes. For DL-based approaches mentioned above, another notable observation is that achieving high robustness requires more than just FDTs—the use of an attack simulator (AS) is essential. This observation prompts us to rethink: Between FDT and AS, which serves as the cornerstone of robustness? Can robust performance be achieved with the AS alone, in the absence of FDT? If FDT proves non-essential, it may be feasible to shift the watermark embedding process from the frequency domain to the temporal domain, potentially enabling fine-grained manipulation of timbre features for improved fidelity.
Figure 1: Two branches of audio watermarking methods. Upper: Frequency-domain watermarking dominates existing SOTA methods. Lower: The proposed temporal-aware watermarking preserves robustness without compromising fine-grained temporal features.
@@
  • Sound Performance. Comprehensive experiments and comparisons with state-of-the-art (SOTA) methods demonstrate the superior fidelity and enhanced robustness.
@@ II-B Extracting Module
Architectures: Gating mechanism[29] has demonstrated its effectiveness in feature extraction for language processing tasks[30] and classification tasks[31]. To fully exploit this advantage, we have designed the gating mechanism to capture temporal domain features, resulting in a Temporal-aware Gated Convolutional Network (TGC), which incorporates the ConvBlock in the extracting module. The TGC follows a dual-branch structure, where the main branch consists of a sequential arrangement of a Conv1d layer, batch normalization, and a sigmoid activation to implement the gating mechanism. In contrast, the shortcut branch contains only a Conv1d layer for acquiring fine-grained features from the input. The outputs of these two branches are merged using the Hadamard Product to produce the final output. This efficient design contributes to the lightweight nature of the ConvBlock.
@@ III Experimental Results and Analysis
This section presents experiments conducted to evaluate the proposed True method based on three primary assessment criteria: fidelity, capacity, and robustness. Additionally, comprehensive experiments and analyses are provided to compare its performance against SOTA methods, further demonstrating its effectiveness.
@@ III-A Experimental Setup
Datasets and Baseline: In the context of voice datasets, other than conventional speech voice datasets like LJSpeech [34], LibriTTS [35], and Aishell3 [36], singing voice datasets like M4Singer [37] and Opencpop [38] are also considered. Moreover, comprehensive comparisons against SOTA watermarking methods are conducted, including handcrafted (HC) methods such as FSVC [16], Normspace [17], and PBML [19], as well as deep learning-based (DL) methods like AudioSeal [21], WavMark [25], TAWM [22], and Groot [33].
@@
Models & Training Settings: In the downsampling block of ICDE, Conv1d layers use a kernel size of 3, a stride of 2, and padding of 2, except for the last layer, which has padding of 1. In the upsampling block, Conv1d layers maintain a kernel size of 3, a stride of 1, and padding of 1, while transposed Conv1d layers have a kernel size of 3, a stride of 2, padding of 2, and output padding of 1, except for the first layer, which uses padding of 1. For the TGCs, both Conv1d layers in the two branches have a kernel size of 3, a stride of 2, and padding of 1. During training, AdamW optimizer [42] is employed with a learning rate of 2e-4. The training is conducted for 40 epochs with a batch size of 16. The hyper-parameters $\lambda_{1}$, $\lambda_{2}$, and $\alpha$ are set to 0.8, 0.1, and 0.3, respectively.
@@ III-C Fidelity and Capacity
Fidelity refers to the imperceptibility of embedded watermarks, measured by the extent to which audio quality is preserved after watermarking. Table II reports both audio quality (for speech and singing voice) and the corresponding watermark extraction accuracy achieved by the proposed True. Additionally, it presents a fidelity comparison against SOTA methods on the LJSpeech dataset. All metrics are computed by comparing the watermarked waveform to its natural counterpart. The results indicate that True achieves excellent speech quality while maintaining reasonable watermark extraction accuracy on speech datasets. Similarly, it demonstrates strong fidelity and extraction performance on singing voice datasets. In comparative experiments with SOTA methods, True ranks first in both PESQ and SSIM metrics, highlighting its superior speech quality.
Capacity which reflects the length of watermarks that can be embedded, is as critical as fidelity in evaluating watermarking performance. Table III illustrates the results of the proposed method on the LJSpeech and Opencpop datasets under varying watermark capacities. Experimental results demonstrate that True is well-suited for high-capacity scenarios, supporting up to 500 bps. On the LJSpeech dataset, the method achieves an average watermark extraction accuracy of 97.84% while maintaining high fidelity, with an average PESQ score of 4.6140. Similarly, for the Opencpop dataset, it achieves an average PESQ of 4.5222 and a recovery accuracy of 98.26%. However, when the capacity exceeds 600 bps, both fidelity and extraction accuracy degrade significantly.
@@ III-D Robustness
We evaluated the robustness of the proposed method across various datasets at a watermark capacity of 100 bps. GN, Pink noise (PN), LP, BP, HP, SPS, ReS, Echo, TS, and Dither were employed for validation. Table IV illustrates the watermark extraction accuracy under these attacks for each dataset. The results show that the proposed method exhibits strong robustness, particularly on singing voice datasets. Specifically, it achieves extraction accuracies of 99.61% and 97.82% under a noise level of 20 dB. Under remanent attacks, True further attains average recovery accuracies of 99.45% and 96.85%, respectively. Although performance on speech datasets is slightly lower than that on singing voice, True still maintains desirable robustness, with average extraction accuracies of 92.33%, 94.85%, and 91.91% across the evaluated conditions.
The proposed True was further compared with SOTA methods on the LJSpeech dataset to evaluate its robustness against various voice post-processing operations, as illustrated in Table V. To account for the capacity limitation of the compared methods, two sets of experiments were conducted at capacities of 32 bps and 100 bps, respectively. At 32 bps, True achieved higher watermark extraction accuracy than all SOTA methods. Notably, it exhibited strong resilience, maintaining an average extraction accuracy of 99.14% even after undergoing six different types of signal attacks. The observed robustness gains can be attributed to ICDE’s deep feature embedding and TGC’s gated decoding, which jointly preserve and extract watermark signals from high-level temporal features. This design makes the method more resilient to typical signal-level distortions. Under the 100 bps configuration, True continued to demonstrate robust performance, particularly in resisting PN and SPS attacks. Although its accuracy under 10 dB GN was slightly lower than that of Groot, True still significantly outperformed the remaining two methods.
@@ IV Conclusion
In this study, we propose True, a temporal-aware robust watermarking method designed to proactively protect the copyrights of diverse waveform types, including both speech and singing voice. To enable seamless watermark embedding, we introduce a content-driven encoder that directly integrates the watermark into the temporal representation of the waveform and reconstructs the watermarked signal end-to-end. For extraction, we develop a temporal-aware gated convolutional network that effectively captures fine-grained features from attacked waveforms, thereby enhancing watermark recovery accuracy. The proposed True surpasses baseline methods in fidelity, supports high-capacity embedding of up to 500 bps, and exhibits strong robustness against a wide range of common waveform distortions.